OpenStudy (genny7):

Need help!! 8 3/10 - 26 1/2

imqwerty (imqwerty):

um can u plese write the question using draw tool

pooja195 (pooja195):

I think they meant this: \[\huge~\rm~ 8\frac{ 3 }{ 10 }-26\frac{ 1 }{ 2 }=? \]

OpenStudy (anonymous):

do you mean \[8\frac{ 3 }{ 10 } - 26\frac{ 1 }{ 2 }\] ?

OpenStudy (genny7):

Yes yes thats what i mean

OpenStudy (anonymous):

so \[\frac{ 83 }{ 10 } - \frac{ 53 }{ 2 }\] = ?

OpenStudy (anonymous):

match denominators and then just subtract

OpenStudy (genny7):

Okay so if i times the denominator by 5 i have to do the same to the top right

OpenStudy (anonymous):

\[\frac{ 83 }{ 10 } - \frac{ -265 }{ 10 }\] is what you are saying you want to do right

OpenStudy (anonymous):

err remove the - on 265 typo

OpenStudy (genny7):

Lol and yes thats what im trying to do

OpenStudy (anonymous):

Yes you are doing it right. from there you can just subtract your values and get your answer

OpenStudy (genny7):

Okay i got -182/10

OpenStudy (genny7):

@Lily2913

OpenStudy (anonymous):

yes that is right but you can simplify it more

OpenStudy (anonymous):

divide by 2 :)

OpenStudy (anonymous):

should i help m8 or yall got this covered

OpenStudy (anonymous):

well time to go to work at best donut place ever dunkin donutsss yahhhhhhhh

OpenStudy (genny7):

So i divided -182 by 2 and i got -91

OpenStudy (genny7):

@Lily2913

OpenStudy (anonymous):

yup -91/5 is the same as -182/10

OpenStudy (genny7):

Oh okay i think i got it from here :) thanks sooooooo much for your help and time! :) @Lily2913

OpenStudy (anonymous):

np. :)